# Evaluating radiation impact on transmon qubits in above and underground facilities

### v1

@@ Abstract
Superconducting qubits can be sensitive to abrupt energy deposits caused by cosmic rays and ambient radioactivity. Previous studies have focused on understanding possible correlated effects over time and distance due to cosmic rays. In this study, for the first time, we directly compare the response of a transmon qubit measured initially at the Fermilab SQMS above-ground facilities and then at the deep underground Gran Sasso Laboratory (INFN-LNGS, Italy). We observe same average qubit lifetime T1 of roughly 80 microseconds at above and underground facilities. We then apply a fast decay detection protocol and investigate the time structure, sensitivity and relative rates of triggered events due to radiation versus intrinsic noise, comparing above and underground performance of several high-coherence qubits. Using gamma sources of variable activity we calibrate the response of the qubit to different levels of radiation in an environment with minimal background radiation. Results indicate that qubits respond to a strong gamma source and it is possible to detect particle impacts. However, when comparing above and underground results, we do not observe a difference in radiation induced-like events for these sapphire and niobium-based transmon qubits. We conclude that the majority of these events are not radiation related and to be attributed to other noise sources which by far dominate single qubit errors in modern transmon qubits.
@@
In recent decades, superconducting circuits have emerged as a leading technology explored in the fields of quantum information [1, 2, 3, 4, 5, 6], photon and particle detectors [7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. In the past years, several research groups have started investigating the effect of particle interactions on the performance of quantum bits (qubits) [17, 18]. There are many studies investigating these potential effects in superconducting quantum devices of various geometries and materials. Some early work pointed to the lifetime of a transmon qubit potentially decreasing when exposed to a strong radioactive source [19]. Moreover, the effect of high energy deposits can rapidly propagate from the initial location of deposition, affecting multiple qubits and potentially leading to a general chip failure [20, 21]. C.D. Wilen et al.[22] observed that the rate of charge jumps experienced by an array of qubits (operated as electrometers) was consistent with the rate expected from interactions due to ambient radioactivity and cosmic rays. Thorbeck et al. [23] did not observe significant radiation effects, but concluded that the impacts due to radioactivity may induce frequency jumps in a dominant loss mechanism, namely two-level systems (TLS) [24, 25, 26, 27, 28, 29, 30, 31], possibly affecting qubit stability.
@@
Other studies have shown that suppressing radioactivity can enhance the performance of superconducting circuits. Operating the device in a radio-pure facility has allowed to significantly improve the locking of a fluxonium qubit at the sweet spot [32] and enhanced the internal quality factor of superconducting resonators [33]. The latter publication also enabled the determination that the substrate on which qubits are deposited, rather than the qubits themselves, is the primary target for particle interactions.
@@
These findings have motivated efforts in further directions aimed at studying and mitigating possible effects of radioactivity. Promising results have been achieved by equipping the chip with “traps” designed to absorb phonons produced (also) by radioactive interactions [34, 35, 36] and using gap engineering [18]. Some groups are also focusing on the components used to construct and operate qubits [37, 38], as it has been demonstrated that some of them, such as the Printed Circuit Boards (PCBs), can produce a sizeable rate of events in the qubit chip [39]. Furthermore, it is also worth mentioning recent proposals of equipping qubits with a device specialised in recognizing particle interactions [40] or with external muon detectors [41, 21, 42].
@@
In this work, for the first time, we study charge-insensitive transmon qubit devices in two laboratories with very distinct radiation environments, as a step of the Superconducting Quantum Materials and Systems (SQMS) Center Round Robin project [39], which aims at studying identical qubits measured at multiple test sites around the world. We first compare average relaxation times measured over several hours, and then employ a fast decay detection protocol to identify reductions in relaxation time at rates faster than 10 kHz and analyze time-bin correlations to determine the characteristics of the events. We study the response of the qubit in a controlled environment, by exposing it to gamma-ray sources of increasing activity. Using the same protocol and analysis strategy, we investigate similar qubits hosted above ground in a typical laboratory environment, to compare effect of radiation on transmon qubits performance and behavior, and estimate timescales at which radiation may become concerning for quantum computational errors.
@@ I Experimental locations
To perform the comparative study, we operated similar transmon chips (charge-insensitive) in two different laboratories: the Quantum Garage of the SQMS Center at Fermilab (FNAL, Illinois, US), and the Ieti Facility located in the deep underground Gran Sasso Laboratory of INFN (INFN-LNGS, L’Aquila, Italy). The chip consisted of a 432 m thick, 7.57.5 mm2 wide HEMEX grade Sapphire substrate. Eight Niobium transmons with different geometries were deposited on the chip (see Supplementary section A); their surface was capped using a 10 nm thick layer of Gold (measured at both sites) or Tantalum (measured at FNAL) to reduce losses due to the formation of Nb2O5 [43]. The capping technique allowed to achieve excellent T1 of many tens and up to hundreds of microseconds, which is important to study the change in qubit’s lifetime at microsecond timescales.
We characterized several qubits of these chips operated at FNAL and at INFN-LNGS, and performed comparisons for either an identical qubit or qubits of similar geometries and materials employed. The detailed parameters of these devices are provided in the Supplementary Section A. Additionally, it is noteworthy that the entire substrate is covered in Niobium creating a ground plane; this may act as a phonon trap, diminishing the phonon propagation and absorption in the active part of the qubit [44, 45].
To predict the rate of impacts due to ambient radiation, we conducted a Monte Carlo simulation using the GEANT4 framework [46] developed in Ref. [39]. We anticipate the total rate of radioactive events interacting with the chip at FNAL and INFN-LNGS to be events/sec and events/sec, respectively. Table 1 provides a summary of the simulation results, with detailed information provided in the Supplementary Section C.
@@
At FNAL (see Fig. 1(a)), the rate is primarily dominated by rays produced by the decay of naturally radioactive isotopes present in the laboratory environment. However, muons generated by cosmic rays also contribute significantly to the overall rate. In contrast, at INFN-LNGS (see Fig. 1(a)), muons are suppressed by six orders of magnitude thanks to the 1.4 km thick rock overburden. Moreover, the radioactivity of the laboratory is substantially reduced by copper and lead shields positioned both inside and around the refrigerator. The last source of radioactivity arises from the materials constituting the qubit chip and its nearby readout components. In a previous study conducted within the SQMS Round Robin project [39], we measured the radioactive content of each component and determined that cables, connectors, amplifiers, circulators, etc., have a negligible impact on the total event rate. The only exception is the PCB, which is typically highly contaminated and located in close proximity to the chip. In this setup, the PCBs are responsible for the whole counting rate reported in the line “Setup” of Table 1. While this rate is negligible at FNAL, it becomes the dominant source of radioactivity at INFN-LNGS, where all external sources of radioactivity are suppressed.
Furthermore, at INFN-LNGS, we utilize Thorium radioactive sources to increase the event rate up to one event every two seconds and investigate the chip behaviour in a “high”-radiation environment.
@@ II Effect of radiation impact
When a radioactive particle impinges on the qubit chip, it releases energy. Muons create long tracks across the chip, whereas rays interact through photo-electric absorption or (mainly via) Compton scattering, producing short-track electrons. In the absence of an electric field, the thousands of charges created along the ionizing track recombine into phonons, which diffuse throughout the chip [22]. As illustrated in Fig. 2(a), when phonons interact with a superconductor, they can break Cooper pairs into quasiparticles. Subsequently, when one or more quasiparticles tunnel through the Josephson Junction, the qubit looses its energy and quickly decays to its ground state [35] with a timescale of 0.16-1.6 s [47, 48]. Since the metallic pads of the transmon host a large number of quasiparticles that take some time to dissipate through normal metal (PCB leads) or lower-energy phonon scattering, these tunneling events can continue to occur, causing the qubit to suffer from strong relaxation for some time. Previous studies claim this low T1 periods to last from one to several tens of milli-seconds [21, 20]. We investigated these type of events in our studies.
We first performed standard T1 experiments on both underground and above ground locations on the identical qubit Q1 and monitored the relaxation times for close to an hour. These experiments were slow and each data point took about 50 seconds to complete. The results, plotted in Fig. 1(b), show similar mean values and fluctuations, typical for transmon qubits [43], demonstrating that radiation has no significant effect at current state-of-the-art transmon qubit lifetimes of 100 microseconds. No sudden drop in T1 values was observed. This was expected as the standard T1 experiments, due to its averaging nature, are not sensitive to fluctuations at milli-second timescales. In order to study possible events at such a time resolution we employed a protocol, tailored for single qubits, utilizing fast decay measurements [20, 18].
@@
Shielded chip | 676.0 | 28.2 |
Shielded chip (2) | – | 124.0 |
44 kBq Th Source | 619.5 | 114.0 |
76 kBq Th Source | 111.7 | 46.3 |
125 kBq Th Source | 60.9 | 28.2 |
161 kBq Th Source | 60.9 | 54.3 |
FNAL Q2 | 234.5 | 104.0 |
FNAL Q3 | 449.7 | 141.9 |
FNAL Q4 | 242.7 | 25.9 |
The fast detection protocol illustrated in Fig. 2(b) begins by resetting the qubit in its excited state . This is accomplished by initially measuring the qubit and applying a conditional -pulse if the qubit is found to be in state via fast feedback. Following a waiting period lasting a few microseconds, the qubit is measured again to determine whether it is in state or . Given a sufficiently long qubit T1, the qubit is highly likely to be found in state . Theoretically, the qubit relaxes to state only a small fraction of the time, given by . However, a radiation event significantly reduces the T1, making it highly probable for the qubit to be detected in state . The sequence is then repeated after another short period, allowing the chip to cool down. Since the effect of quasiparticle generation persists longer than the time taken for one sequence, an event due to radiation typically results in repeated detection in state , as depicted in Fig. 2(c).
In our measurements, we kept the waiting period fixed to 5 s while other parameters like -pulse and readout pulse lengths were similar and reported in the Supplementary Section A. The cooldown period is chosen to ensure that the transmon does not exhibit excessive population in the second excited state . Nevertheless, a simultaneous fit of the , and the second excited state showed that there were many moments in which the (and sometimes even the third excited state ) states were populated. To ensure data consistency, we excluded these time intervals. A thorough investigation of this phenomenon will be mandatory for applications in sensing, as this selection significantly limits the measurement’s live-time (see Table 2).
@@
The readout scheme employed is detailed in the Supplementary Section B. The input signal, filtered through attenuators, low-pass filters, and infrared filters, is sent to the qubit; the output signal is fed into cryogenic triple-junction isolators, and then amplified via a cryogenic HEMT followed by a room temperature amplifier. The qubit readout and control are performed through an RFSoC board using the Quantum Instrumentation Control Kit (QICK) [49]. Each measurement generates a complex number (I/Q values) that are post-processed to distinguish between and based on readout calibration.
@@
We acquired the data for several hours containing several hundred million readouts. During the initial analysis, we simply computed the probability of finding a qubit in for a fixed time window and plot as we moved the window (see Fig. 2(c)). A typical record of as a function of time is shown in Fig. 3(a). The magenta color is the mean of the trace while the fluctuations in the trace look steady. A clear spike around 18.2 seconds corresponds to an event with drastic T1 drop. When zoomed in, these spikes show different pulse shapes as plotted in Fig. 3(b). These pulses generally have a sharp rise but the fall times vary between different instances. As depicted in Fig. 1(b), we found pulses of similar time-profile at both locations irrespective of the specific qubit, pointing to the fact that the drop in relaxation time is not tied to a specific transmon. All pulses observed in our studies last less than ten milli-seconds in agreement with the study in Ref. 21 and thus these pulses require multiple qubit measurements within a milli-second for being detected. For example, the pulses in Fig. 1(c) are visible when a one ms time-window is selected for computing and the signal vanishes when a course time-window of 100 ms is chosen. Clearly, a slow measurement scheme will not be able to resolve individual radiation impact events. In other words, standard T1 measurements, due to its slow nature, will not be affected by radiation impacts (happening at intervals of several seconds) as demonstrated earlier.
@@
A simple detection strategy based on triggering pulses on the data by setting a constant threshold, however, suffers from the problem that a large threshold misses many events, whereas, a lower value picks up a lot of noise arising inherently from the qubit measurement. Furthermore, such a noise is not stable in time leading to unexpected variations when setting a simple constant threshold. As shown in Fig. 3(c), the traces occasionally show noisy behavior which do not follow any adjacent spike. These periods of instability can last from sub-seconds to a minute and will affect regular T1 experiments. Even though it appears to be related to TLS scrambling, we currently do not have a clear understanding of the source of this behavior and it will be a subject of future studies. Consequently, to improve the detection efficiency, we develop a modified strategy discussed below.
III Pulse Reconstruction

### v2

@@ Abstract
Superconducting qubits can be sensitive to abrupt energy deposits caused by cosmic rays and ambient radioactivity. Previous studies have focused on understanding possible correlated effects over time and distance due to cosmic rays. In this study, for the first time, we directly compare the response of a transmon qubit measured initially at the Fermilab SQMS above-ground facilities and then at the deep-underground Gran Sasso Laboratory (INFN-LNGS, Italy). We observe the same average qubit lifetime T1 of roughly 80 microseconds at both above and underground sites despite having very different radiation levels. We then apply a fast decay detection protocol and investigate the time structure, sensitivity and relative rates of triggered events due to radiation versus intrinsic noise, comparing above and underground performance of several high-coherence qubits with similar design. When comparing above and underground results, we do not observe a significant difference in radiation induced-like events for these sapphire and niobium-based transmon qubits. We conclude that the majority of these events are not radiation-related and can be attributed to other noise sources, which predominantly contribute to single qubit errors in modern transmon qubits. Using gamma sources with increasing activity, we assess the response of the qubit to different levels of radiation in an environment with minimal background radiation. Results indicate that qubits respond to gamma sources, demonstrating the possibility to detect particle impacts. However, radiation events can be observed only by significantly reducing other noise sources, which compromises detection efficiency.
@@
In recent decades, superconducting circuits have emerged as a leading technology for applications in the fields of quantum information [1, 2, 3, 4, 5, 6], photon, and particle detectors [7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. In the past years, several research groups have started investigating the effect of particle interactions on the performance of quantum bits (qubits) [17, 18]. There are many studies examining these potential effects in superconducting quantum devices of various geometries and materials. Some early work indicated that the lifetime of a transmon qubit could potentially decrease when exposed to a strong radioactive source [19]. Moreover, the effect of high energy deposits can rapidly propagate from the initial location of deposition, affecting multiple qubits and potentially leading to a general chip failure [20, 21]. C.D. Wilen et al. [22] observed that the rate of charge jumps experienced by an array of charge-sensitive qubits (operated as electrometers) was consistent with the rate expected from interactions due to ambient radioactivity and cosmic rays. Thorbeck et al. [23] did not observe significant radiation effects, but concluded that the impacts due to ionizing radiation may induce frequency jumps in a dominant loss mechanism, namely two-level systems (TLSs) [24, 25, 26, 27, 28, 29, 30, 31], possibly affecting qubit stability.
@@
Other studies have shown that suppressing radioactivity can enhance the performance of superconducting circuits. Operating the device in a radio-pure facility has allowed for the suppression of correlated errors [32] significantly improving the locking of a fluxonium qubit at the sweet spot [33] and enhancing the internal quality factor of superconducting resonators [34]. The latter publication also determined that the substrate on which qubits are deposited, rather than the qubits themselves, is the primary target for particle interactions.
@@
These findings have motivated efforts in further directions aimed at studying and mitigating possible effects of radioactivity. Promising results have been achieved by equipping the chip with “traps” designed to absorb phonons produced (also) by radioactive interactions [35, 36, 37] and using gap engineering [18, 38]. Some groups are also focusing on the components used to construct and operate qubits [39, 40], as it has been demonstrated that some of them, such as the Printed Circuit Boards (PCBs), can produce a sizeable rate of events in the qubit chip [41]. Furthermore, it is also worth mentioning recent proposals of equipping qubits with a device specialized in recognizing particle interactions [42] or with external muon detectors [38, 21, 43].
@@
In this work, for the first time, we study charge-insensitive transmon qubit devices in two laboratories with very distinct radiation environments, as part of the Superconducting Quantum Materials and Systems (SQMS) Center Round Robin project [41], which aims to study identical qubits measured at multiple test sites around the world. We first compare the average relaxation time of a qubit measured over several hours both above ground at Fermilab and deep underground at Gran Sasso Laboratory. We then employ a fast decay detection protocol to investigate radiation-induced events in similar transmon qubits and analyze the time structure, sensitivity, and relative rates of triggered events due to radiation versus intrinsic noise. To further study the qubit’s response to radiation, we expose the superconducting chip to gamma-ray sources with increasing activity in a controlled environment with minimal background radiation and estimate timescales at which radiation may become concerning for quantum computational errors.
@@ I Experimental locations
To perform the comparative study, we operated similar transmon chips (charge-insensitive) in two different laboratories: the Quantum Garage of the SQMS Center at Fermilab (FNAL, Illinois, US), and the Ieti Facility located in the deep-underground Gran Sasso Laboratory of INFN (INFN-LNGS, L’Aquila, Italy). The chip consisted of a 432 m thick, 7.57.5 mm2 wide HEMEX grade Sapphire substrate. It is noteworthy that the entire substrate is covered in Niobium creating a ground plane; this may act as a phonon trap, diminishing the phonon propagation and absorption in the active part of the qubit [44, 45]. Eight Niobium transmons with different geometries were deposited on the chip (see Supplementary section A); their surface was capped using a 10 nm thick layer of Gold (measured at both sites) or Tantalum (measured at FNAL) to reduce losses due to the formation of Nb2O5 [46]. The capping technique allowed us to achieve excellent T1 of many tens and up to hundreds of microseconds, which is crucial for studying changes in a qubit’s lifetime at microsecond timescales. We characterized several qubits of these chips operated at FNAL and at INFN-LNGS, and performed comparisons for either an identical qubit (standard T1 experiment only) or qubits of similar geometries and materials.
To predict the rate of impacts due to ambient radiation, we conducted a Monte Carlo simulation using the GEANT4 framework [47] developed in Ref. [41]. We anticipate the total rate of radiation events interacting within the chip at FNAL and INFN-LNGS to be events/sec and events/sec, respectively. Table 1 provides a summary of the simulation results, with detailed information provided in the Supplementary Section B.
@@
At FNAL (see Fig. 1(a), left panel), the rate is primarily dominated by rays produced by the decay of naturally radioactive isotopes present in the laboratory environment. However, muons generated by cosmic rays also contribute significantly to the overall rate. In contrast, at INFN-LNGS (see Fig. 1(a), right panel), muons are suppressed by six orders of magnitude thanks to the 1.4 km thick rock overburden. Moreover, the radioactivity of the laboratory is substantially reduced by copper and lead shields positioned both inside and around the refrigerator. The last source of radioactivity arises from the materials constituting the qubit chip and its nearby readout components. In a previous study conducted within the SQMS Round Robin project [41], we measured the radioactive content of each component and determined that cables, connectors, amplifiers, circulators, etc., have a negligible impact on the total event rate. The only exception is the PCB, which is typically highly contaminated and located in close proximity to the chip. In this setup, the PCBs are responsible for the whole counting rate reported in the line “Setup” of Table 1. While this rate is negligible at FNAL, it becomes the dominant source of radioactivity at INFN-LNGS, where all external sources of radioactivity are suppressed.
Furthermore, at INFN-LNGS, we utilize calibrated Thorium radioactive sources to increase the event rate up to one event every two seconds and investigate the chip behaviour in a “high”-radiation environment.
@@ II Effect of radiation impact
When a particle impinges on the qubit chip, it releases energy. Muons create long tracks across the chip, whereas rays interact through photo-electric absorption or (mainly via) Compton scattering, producing short-track electrons. In the absence of an electric field, the thousands of charges created along the ionizing track recombine into phonons, which diffuse throughout the chip [22]. As illustrated in Fig. 2(a), when phonons interact with a superconductor, they can break Cooper pairs into quasiparticles. Subsequently, when one or more quasiparticles tunnel through the Josephson Junction, the qubit looses its energy and quickly decays to its ground state [36] with a timescale of 0.16-1.6 s [48, 49]. Since the metallic pads of the transmon host a large number of quasiparticles that take some time to dissipate through normal metal (PCB leads) or lower-energy phonon scattering, these tunneling events can continue to occur, causing the qubit to suffer from strong relaxation for some time. Previous studies claim this low T1 periods last from one to several tens of milliseconds [21, 20]. We investigated these type of events in our studies.
We first performed standard T1 experiments on both underground and above-ground locations on the identical qubit Q1 and monitored the relaxation times for close to an hour. These experiments were slow, and each data point took about 50 seconds to complete. The results, plotted in Fig. 1(b), show similar mean values and fluctuations, typical for transmon qubits [46], demonstrating that radiation has no significant effect at current standard transmon qubit lifetimes of 100 microseconds. No sudden drop in T1 values was observed. This was expected as the standard T1 experiments, due to their averaging nature, are not sensitive to fluctuations at millisecond timescales. In order to study possible events at such a time resolution we employed a protocol, tailored for single qubits, utilizing fast decay measurements [20, 18].
The fast detection protocol illustrated in Fig. 2(b) begins by resetting the qubit in its excited state . This is accomplished by initially measuring the qubit and applying a conditional -pulse if the qubit is found to be in the state via fast feedback. Following a waiting period lasting a few microseconds, the qubit is measured again to determine whether it is in state or . Given a sufficiently long qubit T1, the qubit is highly likely to be found in state . Theoretically, the qubit relaxes to state only a small fraction of the time, given by . However, a radiation event significantly reduces T1, making it highly probable for the qubit to be detected in state . The sequence is then repeated after another short period, allowing the chip to cool down. Since the effect of quasiparticle generation persists longer than the time taken for one sequence, an event due to radiation typically results in repeated detection in state , as depicted in Fig. 2(c).
In our measurements, we select a fixed waiting period of 5 or 10 s; the readout pulse has a length ranging from 3.8 to 8 s (the individual parameters of each measurement are reported in Table S1 of the Supplementary Section A). The -pulse has a negligible length ( 1 s). The cool-down period (10 to 70 s) is chosen to ensure that the transmon does not exhibit excessive population in the second excited state . Nevertheless, a simultaneous fit of the , and the second excited state showed that there were many moments in which the (and sometimes even the third excited state ) states were populated. To ensure data consistency, we exclude these time intervals. A thorough investigation of this phenomenon will be mandatory for applications in sensing, as this selection significantly limits the measurement’s live-time (see Table 2).
@@
Shielded chip | 736.7 | 28.2 |
44 kBq Th Source | 620.7 | 114.0 |
76 kBq Th Source | 124.1 | 46.3 |
125 kBq Th Source | 62.1 | 23.7 |
161 kBq Th Source | 62.1 | 62.1 |
FNAL Q2 | 234.6 | 104.1 |
FNAL Q3 | 449.7 | 92.8 |
FNAL Q4 | 243.1 | 25.9 |
@@
The readout scheme employed is detailed in the Supplementary Section C. The input signal, filtered through attenuators, low-pass filters, and infrared filters, is sent to the qubit; the output signal is fed into cryogenic triple-junction isolators, and then amplified via a cryogenic HEMT followed by a room temperature amplifier. The qubit readout and control are performed through an RFSoC board using the Quantum Instrumentation Control Kit (QICK) [50]. Each measurement generates a complex number (I/Q values) that are post-processed to distinguish between and based on readout calibration.
@@
We acquired the data for several hours containing several hundred million readouts. During the initial analysis, we simply computed the probability of finding a qubit in for a fixed time window and plotted as we moved the window (see Fig. 2(c)). A typical record of as a function of time is shown in Fig. 3(a). The magenta color is the mean of the trace while the fluctuations in the trace look steady. A clear spike around 18.2 seconds corresponds to an event with drastic T1 drop. When zoomed in, these spikes show different pulse shapes as plotted in Fig. 3(b). These pulses generally have a sharp rise, but the fall times vary between different instances. As depicted in Fig. 1(b), we found pulses of similar time-profile at both locations irrespective of the specific qubit, indicating that the drop in relaxation time is not tied to a specific transmon. All pulses observed in our studies last less than ten milliseconds in agreement with the study in Ref. 21, and thus these pulses require multiple qubit measurements within a millisecond for being detected. For example, the pulses in Fig. 1(c) are visible when a one ms time-window is selected for computing and the signal vanishes when a coarse time-window of 100 ms is chosen. Clearly, a slow measurement scheme will not be able to resolve individual radiation impact events. In other words, standard T1 measurements, due to their slow nature, will not be affected by radiation impacts (happening at intervals of several seconds).
@@
A simple detection strategy based on triggering pulses on the data by setting a constant threshold, however, suffers from the problem that a large threshold misses many events, whereas, a lower value picks up a lot of noise arising inherently from the qubit measurement. Furthermore, such a noise is not stable in time leading to unexpected variations in the rate. As shown in Fig. 3(c), the traces occasionally show noisy behavior which do not follow any adjacent spike. These periods of instability can last from sub-seconds to a minute and will affect regular T1 experiments. Even though it appears to be related to TLS scrambling, we currently do not have a clear understanding of the source of this behavior, and it will be a subject of future studies. To better identify radiation-induced events, we developed a modified strategy discussed below.
III Event Selection